OpenStudy (anonymous):

What multiplies to -255, but adds to 0? I've been trying to figure this out for a while, and can't figure it out.

OpenStudy (anonymous):

zero.....not really sure

OpenStudy (anonymous):

can you rewrite the ques?

OpenStudy (johnweldon1993):

i assume your factoring? lol well if the two numbers add to 0...they have to be the same number right? so what is 255 / 2? not a whole number is it? so there is no solution this is if it's factoring of course

OpenStudy (anonymous):

\[ 16\times -16 = -256 \]So it's pretty close to \(16\).... \[ x\times x = x^2 = -255\\ x+x = 2x = 0 \]

OpenStudy (anonymous):

Or are you asking for a pair of numbers?